Let X be a random variable that is normally distributed with a mean of 10 and a

standard deviation of 2. Assume that a sample of 25 observations are taken from

this distribution, What is the interquartile range of the distribution of the sample

means.


1
Expert's answer
2022-03-15T05:17:08-0400

Q1= "\mu" - (0.675)*"\frac{\sigma}{\sqrt{n}}"

Q3="\mu" + (0.675)*"\frac{\sigma}{\sqrt{n}}"

Given "\mu" = 10, "\sigma" =2 and n=25.

Then,

Q1="10-(0.675)\frac{2}{\sqrt{25}}" =9.73

Q3= "10+(0.675)\frac{2}{\sqrt{25}}" =10.27

The interquartile range (IQR) = Q3-Q1.

=10.27 - 9.73

=0.54

Answer: 0.54